the sum of 3 numbers is 84. the first number is 2 times the second number and the second number is 2 times the third number. what is the value of each number
Answer by mananth(16946)

third number = x
second number = 2x
first number = 2*2x=4x
x+2x+4x=847x=84
/7
x=12
12,24, 48